what does 5! mean ?

Grade:12

35 Answers

Prajwal Kavad
8 years ago
5! means 5 factorial, it si equal to 5x4x3x2x1=120
mamam
8 years ago
thank you
V. Harshit -
8 years ago
5!=120
Kalyani Jayachandran Menon
8 years ago
In Mathematics, ! stands for factorial.
 
For example, x! means the product of all natural numbers till x(including x).
 
Hope it helps..
Adarsh
8 years ago
5! stands for 5 factorial and its value is 5*4*3*2*1=120
KARTHIK
8 years ago
120
venky
8 years ago
5! stands for 5 factorial and its value is5!=5*4*3*2*1
pa1
8 years ago
5! stands for 5 factorial and its value is 5*4*3*2*1=120
Gaddam Chethan
8 years ago
5! means 5*4*3*2*1. in same way 8! means 8*7*6*5*4*3*2*1*. for any number you have to define like this only                         
Shreyas Patil
7 years ago
120.
utpal raj
7 years ago
it is
YASWANTH
7 years ago
5! means multiplying the possible numericcals from the given number degrading to 1..
5*4*3*2*1... so that 5*4*3*2*1 gives 120
 
venkata ramana
7 years ago
hi friend5! stands for 5 factorial
so the answer is 5*4*3*2*1=120
all the best.
YASWANTH KISHOR
7 years ago
5! means multiplying the possible factors from 5 to degrading to 1 tha is 5*4*3*2*1= 120 and all the best be first in ur future
Ritika Das
6 years ago
5! means factorial, which means we have to multiply the predecessors of 5 only natural numbers up to 1. Therefore we can get the result as 5*4*3*2*1, which is equal to 120.
Agastya Varma
6 years ago
! Is the symbol for factorial. Factorial is the product of all the natural numbers less that the number given including the number. 5! = 1*2*3*4*5= 120.
 
 
 
saiteja Karingu
6 years ago
Dear Students,
Lets assume A be a digit then the notation A! implies the product of all the number before A From 1 to A .
For example :  1! = 1 , 2! = 2 × 1 = 2 , 3! = 3 × 2 × 1 = 6, 4! = 4 × 3 × 2 × 1 = 24 
Now 5! = 5 × 4 × 3 × 2 × 1 = 20 × 6 = 120
Remember that 0! is "1". You get why it is in chapter called permuations in class 11th.
Happy learning ,
All the best.
 
 
Sriesh
6 years ago
5! Is written as 5 factorialIt is 1*2*3*4*5Another example-7!1*2*3*4*5*6*7This means that u have to multiply everything before the number
swarup kumar majee
6 years ago
it is 5 factorial.5 factorial means 5*4*3*2*1=120
so,the value of 5! is=120
this means you have to multiply that number and all numbers before that leaving 0.
Harshit Panwar
6 years ago
5factorial means 5×4×3×2×1
5!=5×4×3×2×1=120
Hope it helps you...........................................,....
ankan sen
6 years ago
5! means 5 factorial
it is calculated as 5*4*3*2*1=120
this is in the chapter permutation and combination
 
Abhinav Gupta
6 years ago
 
 
Dear Students,
Lets assume A be a digit then the notation A! implies the product of all the number before A From 1 to A .
For example :  1! = 1 , 2! = 2 × 1 = 2 , 3! = 3 × 2 × 1 = 6, 4! = 4 × 3 × 2 × 1 = 24 
Now 5! = 5 × 4 × 3 × 2 × 1 = 20 × 6 = 120
Vivek
6 years ago
5! =120 Because 5! Represents 5×4×3×2×1=120 Factorial of any no. Can be found by multilpyng no. by no. One less than it till it doent goes to 1
praveen kumar
6 years ago
5! means the function of factorial it means the value of 5!=5*4*3*2*1=120.
10!=10*9*8*7*6*5*4*3*2*1 and other numbers as same.
thank you.
unnati kasaundhan
6 years ago
5! Is says to be factorial five 5! = 5*4*3*2*1if the given question is given as.................n! then General formula to solve this ........n! = n*(n-1)*(n-2)*(n-3)*(n-4)*................*(n-n)(n-n)!=0!=1As......4!=4*(4-1)*(4-2)*(4-3)*(4-4) =4*3*2 *1*0! =24*0! =24
Priyansh
6 years ago
5! is read as 5 factorial ans has a value 5x4x3x2x1 that is 120.........................................................................................................
Ashna
6 years ago
5!=1×2×3×4×5=120 it is read as factorial 5.Its very easy to write factorials.You can try it with other numbers too.
Ashna
6 years ago
Factorial is a term which you will study in class 11 in a chapter named as permutation and combination so you don`t need to worry now .But value of 5!=120
Akachdeep Medhi
6 years ago
5! or L5 means 5 factorial.
Mathematically,
             5! = 5 X 4 X 3 X 2 X 1
                 = 20 X 6 X 1
                 = 120
With regards Akash.
sahil
6 years ago
Factorial is a term which you will study in class 11 but if you wanna know then 5! Means 5×4×3×2×1 you will study this in higher class and it is very useful
Varun
6 years ago
5! Means 5*4*3*2*1=120 It means that is a factorial of 5..........................................................
Siddharth Singh
6 years ago
5!= 5*4*3*2*1=120.....read as five factorial...... mathematical operation......Thanksgiving...............................
Soma Mesh
5 years ago
It means factorial of that no . it's according to Pascal's triangle . 5!= 5*4*3*2*1 which giver 110+10
Sai Soumya
5 years ago
Dear Student . 
5! , known as 5 factorial means 
5! = 5*4*3*2*1=120
similarly all other factorials are counted in the same way . 
john cena
5 years ago
this means 5 factorial
the term factorial means product of n natural numbers
if i would say n! so it means that product of natural numbers upto n
so 5! means 1*2*3*4*5
which is equal to 120